See Figure 1 for an overview of the Simplicissimus architecture.

Figure 1: A user expression, translated into Gnu’s abstract syntax tree (upper box), mapped to the internal representation in Simplicissimus, rewritten by the core engine (lower box), and mapped back.

In Figure 2 we have the same diagram but it’s reconfigured a bit.

Figure 2: Same diagram as in Figure 1, but drawn a little larger, stretched in one dimension, and rotated 90 degrees counterclockwise.